OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#9122 — Improvement Lille 10G optic loop
Scheduled Maintenance Report for Web Cloud
Completed
We are going to insert a new Lambda into the fibre optic to increase the capacity on our NRAs.
We are working on this loop: PIL59 - BLE59 - BOI59 - FAC59 - SEC59 - LIB59.
No impact on traffic is expected. Works will commence Friday August 9th at 14:00.

Update(s):

Date: 2013-08-16 16:08:53 UTC
We have found the issue on the securisation side, the
optic levels were too weak. We will intervene on Monday
August 19th to insert an EDFA amp, in order to obtain
the correct optic power.

Date: 2013-08-16 11:07:46 UTC
We have succeeded to raise one side of the 10G loop,
however, we are still having reception issues with securisation.
We are going to intervene this afternoon in order to
perform tests on the fibre optic.
We will start the intervention at 3pm Friday August 16th.

Date: 2013-08-12 15:12:57 UTC
We are moving the amp to boi59, securisation is operational in case of power failure.

Date: 2013-08-12 14:23:57 UTC
We are continuing with the intervention on the Lille optic loop. We
are going to reconfigure the EDFA amps in order to move the loop
onto 10G.

Date: 2013-08-09 16:47:24 UTC
We have an optic attenuation issue. We are halting the intervention. Optimisation works are required on the optic loop. We will start again on Monday 12th.

Securisation is operational in case of power failure.

Date: 2013-08-09 13:55:21 UTC
We are working on the first side of the loop.

Date: 2013-08-09 13:54:35 UTC
We are starting the maintenance works, we are inserting the 10G cards on the relevant sites.
Posted Aug 09, 2013 - 10:44 UTC